# Environment-Driven Online LiDAR-Camera Extrinsic Calibration

### v2

@@ IV-B2 Evaluation on MIAS-LCEC Dataset
IV-C Ablation Study and Analysis
| Consistency | Left Camera | Right Camera | |||
| Structural | Textural | (∘) | (m) | (∘) | (m) |
| 0.360 | 0.114 | 0.476 | 0.185 | ||
| ✓ | 0.340 | 0.102 | 0.473 | 0.152 | |
| ✓ | 0.328 | 0.109 | 0.435 | 0.146 | |
| ✓ | ✓ | 0.295 | 0.082 | 0.336 | 0.118 |
| Components | Left Camera | Right Camera | ||||||||||||||||
| Multi-View | Multi Scene | Magnitude | Rotation Error (∘) | Translation Error (m) | Magnitude | Rotation Error (∘) | Translation Error (m) | |||||||||||
| Intensity | Depth | (∘) | (m) | Yaw | Pitch | Roll | X | Y | Z | (∘) | (m) | Yaw | Pitch | Roll | X | Y | Z | |
| 1.625 | 0.459 | 0.820 | 0.669 | 0.899 | 0.247 | 0.235 | 0.211 | 1.620 | 0.472 | 0.915 | 0.691 | 0.807 | 0.257 | 0.254 | 0.197 | |||
| ✓ | 1.387 | 0.358 | 0.755 | 0.579 | 0.711 | 0.189 | 0.208 | 0.153 | 1.641 | 0.459 | 1.012 | 0.679 | 0.738 | 0.257 | 0.257 | 0.178 | ||
| ✓ | ✓ | 1.125 | 0.280 | 0.534 | 0.534 | 0.613 | 0.134 | 0.151 | 0.136 | 1.425 | 0.354 | 0.856 | 0.563 | 0.679 | 0.186 | 0.205 | 0.138 | |
| ✓ | 0.406 | 0.155 | 0.180 | 0.201 | 0.223 | 0.119 | 0.067 | 0.049 | 0.447 | 0.192 | 0.227 | 0.243 | 0.211 | 0.148 | 0.094 | 0.051 | ||
| ✓ | ✓ | 0.339 | 0.110 | 0.179 | 0.167 | 0.162 | 0.069 | 0.063 | 0.039 | 0.480 | 0.138 | 0.322 | 0.239 | 0.150 | 0.096 | 0.084 | 0.033 | |
| ✓ | ✓ | ✓ | 0.295 | 0.082 | 0.117 | 0.176 | 0.150 | 0.051 | 0.044 | 0.032 | 0.336 | 0.118 | 0.216 | 0.168 | 0.121 | 0.083 | 0.067 | 0.032 |
To validate the contribution of structural consistency and textural consistency in DPCM, we conducted an ablation study comparing their individual and combined effects. As shown in Table V, both components significantly improve calibration accuracy. Specifically, structural consistency preserves local geometric features through mask-guided alignment, while textural consistency evaluates visual similarity around matched correspondences. Their combination enables DPCM to maintain robust matching performance even in challenging scenarios with sparse or occluded point clouds, which is a capability that existing methods (DVL and MIAS-LCEC) struggle to achieve. Additionally, MIAS-LCEC only compares the corner points within the matched segmented mask, while EdO-LCEC not only considers the corner points of the matched mask but also includes those of other unmatched masks. This allows EdO-LCEC to achieve global attention when searching the potential available matches when LiDAR frames are sparse.
To explore the contribution of the generalizable scene discriminator as well as the multi-view and multi-scene joint optimization, we evaluate our algorithm’s performance with and without multi-view (including both textural and spatial perspective views) and multi-scene optimizations on the 00 sequence of KITTI odometry. The results presented in Table VI demonstrate that lacking any of these components significantly degrades calibration performance. In particular, calibration errors increase substantially when only single-view calibration is applied, as it lacks the comprehensive constraints provided by multi-view inputs. Additionally, the joint optimization from multiple scenes significantly improved the calibration accuracy compared to those only under multi-view optimization. These results confirm the advantage of incorporating spatial and textural constraints from multiple views and scenes, validating the robustness and adaptability of our environment-driven calibration strategy.
@@ VI Details of the Algorithm Implementation
[…]
to represent this affine transformation (mentioned in Sect. III-B), where represents the rotation matrix, denotes the translation vector, and represents the scaling factor. We can assume that after the affine transformation, any points within a given mask in the LIP image perfectly align with the corresponding points from the RGB image, and thus:
where and represent the updated coordinates of the instance’s center and corner points in the virtual image, respectively. In this case, can be obtained using the following expression:
where
is the angle between the vectors originating from the mask centers and pointing to their respective matched corner points. can then be expressed as follows:
which represents the ratio between the areas of the bounding boxes associated with the RGB image and the LIP image. Finally, according to (13), can be obtained as follows:
In DPCM, we utilize the affine transformation estimated by (14)-(17) as a semantic prior to guide the dense correspondence matching between LiDAR and camera.
Additionally, the component of the structural consistency metioned in Sect. III-B is defined as follows:
which represents the similarity of the neighboring vertices between the current and target corner points.
For the spatial-temporal relative pose optimization, as presented in Sect. III-C, we utilize a Gaussian kernel to calculate the depth-normalized reprojection error. In detail, is defined as follows:
where is the true average depth of the correspondences.
@@
VII Calibration Parameters
Calibration Parameters for Left RGB Camera
Calibration Parameters for Right RGB Camera
An example of the data fusion result using our extrinsic calibration parameters for the left camera in KITTI odometry is shown in Fig. 9.
VIII Details of Experimental Configuration
@@
| RegNet [43] | IV 2017 | Learning-based | CNN | Medium | Medium | No |
| CalibNet [25] | IROS 2018 | Learning-based | Geometrically Supervised Deep Learning Network | High | Low | Yes |
@@
| CalibRCNN [23] | IROS 2020 | Learning-based | RCNN | Medium | Medium | No |
| CalibDNN [24] | SPIE 2021 | Learning-based | Deep Learning Network | Medium | Medium | No |
| LCCNet [22] | CVPR 2021 | Learning-based | CNN, L1-Loss | High | Medium | Yes |
@@
VIII-A Datasets
@@
We have conducted extensive experiments on KITTI odometry and MIAS-LCEC (including MIAS-LCEC-TF70 and MIAS-LCEC-TF360) datasets. Now we provide their details:
KITTI odometry: KITTI odometry is a large-scale public dataset recorded using a vehicle equipped with two RGB cameras Point Grey Flea 2 and one LiDAR of type Velodyne HDL-64E. Sensor data is captured at 10 Hz. We utilize the first 10 sequences (00-09) for evaluation.
MIAS-LCEC-TF70: MIAS-LCEC-TF70 is a diverse and challenging dataset that contains 60 pairs of 4D point clouds (including spatial coordinates with intensity data) and RGB images, collected using a Livox Mid-70 LiDAR and a MindVision SUA202GC camera, from a variety of indoor and outdoor environments, under various scenarios as well as different weather and illumination conditions. This dataset is divided into six subsets: residential community, urban freeway, building, challenging weather, indoor, and challenging illumination, to comprehensively evaluate the algorithm performance.
MIAS-LCEC-TF360: MIAS-LCEC-TF360 contains 12 pairs of 4D point clouds and RGB images, collected using a Livox Mid-360 LiDAR and a MindVision SUA202GC camera from both indoor and outdoor environments. Since the Livox Mid-360 LiDAR has a scanning range of 360∘, it produces a sparser point cloud compared to that generated by the Livox Mid-70 LiDAR. Additionally, the significant difference in the FoV between this type of LiDAR and the camera results in only a small overlap in the collected data. Consequently, this dataset is particularly well-suited for evaluating the adaptability of algorithms to challenging scenarios.
VIII-B Data Preparation
@@
The dataset statistics of KITTI odometry are presented in Table VIII. It can be observed that there are a total of 24500 pairs of point clouds and images in KITTI odometry. Note that most of the SoTA comparison methods are slow to complete calibration with a single frame. For the comparison algorithms that provided source code, we performed calibration every ten frames (except for 2 frames in 04 sequence), enhancing the efficiency of the experiments without compromising the validity of the evaluation.
The data preparation pipeline for our MIAS-LCEC datasets is illustrated in Fig. 11. Our experimental platform integrates two DJI Livox solid-state LiDARs (Mid-70 and Mid-360) and a MindVision MV-SUA202GC global-shutter CMOS camera for synchronized cross-modal data acquisition. Both LiDARs operate at 200,000 points/s. However, the Mid-70 captures dual returns while the Mid-360 records only the first return. For each scene, we accumulate 15-20 seconds of LiDAR scans to ensure dense point cloud coverage, paired with RGB images at 1,200800 pixel resolution captured by the camera, whose intrinsic parameters were pre-calibrated offline and remained fixed throughout all experiments.
VIII-C Details of Comparison Methods
The details of the comparison methods are illustrated in Table VII. For algorithms that did not release their code, we directly reference the results published in their papers. Notably, RegNet [43], Borer et al. [21], and CalibDNN [24] employ different experimental datasets and evaluation processes, which limits the potential for a fully quantitative comparison. Furthermore, these methods do not specify which frames from the KITTI dataset were used for training and testing. Therefore, we rely on the published results to compare these approaches to others on the KITTI odometry 00 sequence.

### v3

@@ IV-A Experimental Setup and Evaluation Metrics
| Approach | Magnitude | Rotation Error (∘) | Translation Error (m) | |||||
| (∘) | (m) | Yaw | Pitch | Roll | X | Y | Z | |
| CRLF [40] | 2.446 | 14.82 | 0.537 | 0.454 | 2.317 | 12.25 | 4.159 | 4.852 |
| UMich [20] | 4.340 | 0.390 | 0.524 | 2.118 | 3.185 | 0.297 | 0.071 | 0.200 |
| HKU-Mars [19] | 36.69 | 17.55 | 16.150 | 24.91 | 19.69 | 11.216 | 5.878 | 7.171 |
| DVL [32] | 108.9 | 3.371 | 42.50 | 71.71 | 70.88 | 1.672 | 1.874 | 1.380 |
| MIAS-LCEC [4] | 2.731 | 0.688 | 0.625 | 0.463 | 2.570 | 0.554 | 0.072 | 0.325 |
| EdO-LCEC (Ours) | 1.090 | 0.261 | 0.414 | 0.356 | 0.842 | 0.202 | 0.076 | 0.100 |
IV-B Comparison with State-of-the-Art Method
In this section, quantitative comparisons with SoTA approaches on three datasets are presented in Fig. 6, Tables I 222The reproduced results of LCCNet yield higher errors compared to those reported in their paper., II, III, and IV. Additionally, qualitative results are illustrated in Figs. 5 and 7.
IV-B1 Evaluation on KITTI Odometry
The results shown in Table I and II suggest that, with the exception of sequences 01 and 04, our method achieves SoTA performance across the ten sequences (00-09) in KITTI odometry. Specifically, in the 00 sequence, EdO-LCEC reduces the by around 35.2-99.8% and the by 16.1-98.8% for the left camera, and reduces the by around 46.9-99.7% and the by 13.9-97.6% for the right camera. Additionally, according to Fig. 7, it can be observed that the point cloud of a single frame in KITTI is so sparse that the other approaches behave poorly. In contrast, our proposed method overcomes this difficulty and achieves high-quality data fusion through the calibration result. We attribute these performance improvements to our multi-view and multi-scene joint optimization. Merging the optimal matching results from multiple views and scenes maximizes the number of reliable correspondences and ultimately improves overall calibration accuracy.
IV-B2 Evaluation on MIAS-LCEC Dataset
@@
IV-B3 Comparison with SoTA Approaches on nuScenes
Following the official split of nuScenes, we adopt the 150 test scenes in the v1.0-test set to construct the evaluation data. The experimental results, summarized in Table V, show that EdO-LCEC significantly outperforms prior methods. We attribute this improvement to the robust correspondences established by DPCM. Although existing approaches obtain dense correspondences from accumulated point clouds captured by solid-state LiDARs, the accuracy of these correspondences is limited and does not generalize well to unseen scenarios. The substantial modality gap between LiDAR point clouds and camera images further hinders prior methods from achieving high pose estimation accuracy. Moreover, when applied to sparse point clouds captured by mechanical LiDARs, these methods suffer a severe degradation in performance, largely due to the pronounced differences in feature density between LiDAR and camera data. In contrast, our approach leverages cross-modal segmented masks to bridge the modality gap. The corner points extracted from these masks inherently encode reliable semantic information, which is critical for effective correspondence matching. DPCM not only increases the number of correspondences but also improves their geometric distribution, making them better suited for accurate pose estimation.
IV-C Ablation Study and Analysis
Left Camera | Right Camera | ||||
| Structural | Textural | (∘) | (m) | (∘) | (m) | |
| 2.222 | 0.792 | 2.600 | 0.896 | |||
| ✓ | 1.227 | 0.304 | 1.533 | 0.411 | ||
| ✓ | 2.416 | 0.827 | 2.633 | 0.984 | ||
| ✓ | ✓ | 1.125 | 0.278 | 1.425 | 0.354 | |
| Components | Left Camera | Right Camera | ||||||||||||||||||
| Multi-View |
Magnitude | Rotation Error (∘) | Translation Error (m) | Magnitude | Rotation Error (∘) | Translation Error (m) | |||||||||||||
| Intensity | Depth | (∘) | (m) | Yaw | Pitch | Roll | X | Y | Z | (∘) | (m) | Yaw | Pitch | Roll | X | Y | Z | |||
| 1.625 | 0.457 | 0.820 | 0.669 | 0.899 | 0.247 | 0.232 | 0.211 | 1.620 | 0.472 | 0.915 | 0.691 | 0.807 | 0.257 | 0.254 | 0.197 | |||||
| ✓ | 1.387 | 0.357 | 0.755 | 0.579 | 0.711 | 0.189 | 0.205 | 0.152 | 1.641 | 0.459 | 1.012 | 0.679 | 0.738 | 0.257 | 0.257 | 0.178 | ||||
| ✓ | ✓ | 1.125 | 0.278 | 0.534 | 0.534 | 0.613 | 0.134 | 0.148 | 0.136 | 1.425 | 0.354 | 0.856 | 0.563 | 0.679 | 0.186 | 0.205 | 0.138 | |||
| ✓ | 0.406 | 0.151 | 0.180 | 0.201 | 0.223 | 0.119 | 0.058 | 0.049 | 0.447 | 0.192 | 0.227 | 0.243 | 0.211 | 0.148 | 0.094 | 0.051 | ||||
| ✓ | ✓ | 0.339 | 0.106 | 0.179 | 0.167 | 0.162 | 0.069 | 0.056 | 0.039 | 0.480 | 0.138 | 0.322 | 0.239 | 0.150 | 0.096 | 0.084 | 0.033 | |||
| ✓ | ✓ | ✓ | 0.295 | 0.078 | 0.117 | 0.176 | 0.150 | 0.051 | 0.038 | 0.032 | 0.336 | 0.118 | 0.216 | 0.168 | 0.121 | 0.083 | 0.067 | 0.032 | ||
| Components | Camera 1 | Camera 2 | ||||||
| Multi-View |
Magnitude | Magnitude | |||||
| Intensity | Depth | (∘) | (m) | (∘) | (m) | |||
| 2.232 | 0.524 | 1.909 | 0.440 | |||||
| ✓ | 1.832 | 0.402 | 1.882 | 0.415 | ||||
| ✓ | ✓ | 1.496 | 0.323 | 1.506 | 0.311 | |||
| ✓ | 0.914 | 0.197 | 0.686 | 0.134 | ||||
| ✓ | ✓ | 0.755 | 0.135 | 0.612 | 0.085 | |||
| ✓ | ✓ | ✓ | 0.605 | 0.079 | 0.498 | 0.061 | ||
@@ IV-C Ablation Study and Analysis
To validate the contribution of structural consistency and textural consistency in DPCM, we conducted an ablation study comparing their individual and combined effects. Multi-scene optimization is not used in this ablation study to better demonstrate the influence of the two consistencies. As shown in Table VI, when both are used, the two components significantly improve calibration accuracy. Specifically, structural consistency preserves local geometric features through mask-guided alignment, while textural consistency evaluates visual similarity around matched correspondences. Their combination enables DPCM to maintain robust matching performance even in challenging scenarios with sparse or occluded point clouds, which is a capability that existing methods (such as DVL and MIAS-LCEC) struggle to achieve. Additionally, MIAS-LCEC only compares the corner points within the matched segmented mask, while EdO-LCEC not only considers the corner points of the matched mask but also includes those of other unmatched masks. This allows EdO-LCEC to achieve global attention when searching for potential available matches when LiDAR frames are sparse.
Furthermore, we explore the contribution of the generalizable scene discriminator as well as the multi-view and multi-scene joint optimization. In this ablation study, the algorithm’s performance is comprehensively evaluated on the 00 sequence of KITTI odometry and KITTI-360, both with and without multi-view optimization (including both textural and spatial perspective views) and multi-scene optimization. The results presented in Table VII and VIII demonstrates that lacking any of these components significantly degrades calibration performance. In particular, calibration errors increase substantially when only single-view calibration is applied, as it lacks the comprehensive constraints provided by multi-view inputs. Additionally, the joint optimization from multiple scenes significantly improved the calibration accuracy compared to that under only multi-view optimization. These results confirm the advantage of incorporating spatial and textural constraints from multiple views and scenes, validating the robustness and adaptability of our environment-driven calibration strategy.
@@ VI Details of the Algorithm Implementation
VI-A Estimation of the 4-DoF Similarity Transformation
As presented in the “Dual-Path Correspondence Matching” in Sect. III-B, we initially adopted the method described in [4] for cross-modal mask matching. [4] proved that when two 3D LiDAR points (reference) and (target) in the virtual camera coordinate system are close in depth, the relationship of their projection coordinates and on virtual image and camera image can be expressed by a following 4-DoF similarity transformation:
where represents the rotation matrix, denotes the translation vector, and represents the scaling factor. We can assume that after the similarity transformation, any points within a given mask in the LIP image perfectly align with the corresponding points from the RGB image, and thus:
where and represent the updated coordinates of the instance’s center and corner points in the virtual image, respectively. In DPCM, we adopt the method proposed in [4] to estimate the similarity transformation, which serves as a semantic prior to guide dense correspondence matching between the LiDAR and camera modalities. When the calibration scenarios become complex or the initial misalignment between the LiDAR and camera is relatively large, the estimation of the rotational-scale component in the similarity transformation may become unstable. In such real-world cases, can be approximated by an identity matrix, allowing the transformation process to primarily depend on the translation component .
VI-B Components in the Structural Consistency
In the algorithm implementation, the structural consistency component introduced in Sect. III-B can be defined as follows:
The function in (13) measures the similarity between a neighboring vertex of the current corner point and that of the corresponding target corner point, and is formulated as:
where is a constant scale parameter.
VI-C Components in the Multi-View Optimization
For the multi-view optimization, as presented in Sect. III-C, we utilize a Gaussian kernel to calculate the depth-normalized reprojection error. In detail, is defined as follows:
where is the true average depth of the correspondences. The normalized depth in the equation is the normalization result of the point cloud depth in the range . The calculation of is defined as follows:
where and are the minimum and maximum depth values of the input point cloud in the virtual camera perspective.
@@ VI-C Components in the Multi-View Optimization
VI-D Configuration of Multi-Scene Optimization
As shown in Table VII and VIII in Sect. IV-C, multi-scene optimization significantly improves calibration accuracy compared to multi-view optimization. Using more frames leads to greater accuracy improvements from multi-scene optimization. In our experiments, we used 20 frames to obtain the results reported in the table. For real-world applications, we recommend using at least 5 frames to achieve stable multi-scene calibration. When the point cloud is relatively dense, as in the TF70 dataset, a single frame is enough to produce accurate results by using the multi-view optimization.
VI-E Discussion on Virtual Camera Generation
As a crucial component of the scene discriminator, the virtual camera generation process ensures proper positioning for capturing LiDAR projections. The primary goal of virtual camera generation is to maximize the capture of matchable features for correspondences between LiDAR point clouds and camera images. Therefore, virtual cameras must be judiciously positioned and deployed in sufficient numbers. EdO-LCEC estimates the feature density of the environment, derived from LiDAR projections and camera images, to determine the number of virtual cameras required to balance calibration coverage and computational cost. The virtual cameras are arranged along the , , and axes on a sphere centered at the LiDAR perspective origin. This adaptive generation approach allows dynamic adjustment of the virtual camera count according to environmental complexity, making it suitable for diverse and uncertain operating conditions. For instance, in the relatively sparse KITTI dataset, more virtual cameras (around 7) may be needed to ensure adequate coverage, whereas in dense scenarios such as the TF70 dataset, fewer virtual cameras (about 1-2) suffice, avoiding unnecessary computational overhead. Extensive experiments in Sect. IV demonstrates that this feature-density-based generation strategy performs robustly across challenging scenarios and exhibits strong environmental adaptivity. We contend that this virtual camera generation method is optimal within our proposed environment-driven calibration framework.
@@ VI-E Discussion on Virtual Camera Generation
| RegNet [44] | IV 2017 | Learning-based | CNN, Direct Pose Regression | Medium | Medium | No |
| CalibNet [28] | IROS 2018 | Learning-based | Geometrically Supervised DNN | High | Low | Yes |
@@
| CalibRCNN [26] | IROS 2020 | Learning-based | DCNN, LSTM, Direct Pose Regression | Medium | Medium | No |
| CalibDNN [27] | SPIE 2021 | Learning-based | CNN, Direct Pose Regression | Medium | Medium | No |
| LCCNet [25] | CVPR 2021 | Learning-based | CNN, Cost Volume, Direct Pose Regression | High | Medium | Yes |
@@
In our preliminary experiments, we attempted to fix the number of virtual cameras according to the complexity of the calibration data. In relatively simple environments with few dynamic objects, such as a controlled calibration room, this method proved stable and easy to configure, given available computational resources. While this manual strategy enhanced calibration performance in certain specific cases, it lacked generalizability across more complex and diverse scenarios. The required parameters vary with environmental changes, and the absence of environmental perception prevents the algorithm from automatically adapting to unseen and challenging conditions.
Apart from this manual way to generate virtual cameras, we have also considered two possible alternatives that can achieve automatic generation. Here, we present these strategies and provide tailored application recommendations for different scenarios:
One of the automatic methods to determine the number of virtual cameras is to compute the ratio of LiDAR to camera fields of view, which allows deducing the minimum number of virtual cameras required. As depicted in Fig. 10. (a), each virtual camera matches the intrinsic FoV of the real camera, and collectively they cover the entire LiDAR scan range. A straightforward formula for determining the required number of virtual cameras is:
(19) where is the LiDAR’s field of view, is the camera’s field of view, and denotes the ceiling function to ensure full coverage of the LiDAR scan. The parameter defines the virtual camera density when the FoV ratio between the LiDAR and camera is equal to 1. If the FoVs of LiDAR and camera are unknown. We can estimate them using the following method:
For the FoV of LiDAR, we can obtain an estimate of the FoV using the captured point clouds. Let the LiDAR point cloud be in the sensor frame. For each point define the horizontal angle and vertical angle by the following equation:
(20) The horizontal and vertical fields of view can then be obtained as follows:
(21) For the camera, its FoV can be computed as follows:
(22) where is the relevant sensor dimension (e.g., width or height) and is the focal length. Using these quantities, one can decide virtual camera counts via:
(23) with the choice of horizontal or vertical FoV depending on the coverage requirement. This FoV-based approach is particularly effective when the FoV difference between the LiDAR and camera is moderate. If the FoV gap is too large, a substantial portion of the projection may remain unused for feature matching. Conversely, if the FoV gap is too small, too few projections are produced, which can cause failures in scenarios with sparse LiDAR point clouds. While this strategy overcomes the feature matching difficulties introduced by FoV discrepancies, it does not address the modality gap or the differences in feature density between LiDAR and camera. Moreover, the absence of an environmental perception mechanism limits its adaptability, making it unsuitable for an environment-driven calibration framework.
If an initial estimate of the extrinsic parameters is available, the number of virtual cameras can also be determined by the distance between the LiDAR origin and the initial estimate. As illustrated in Fig. 10. (b), given virtual cameras per unit distance, the number of virtual cameras can be calculated as
(24) and their positions are distributed along the translation vector from the LiDAR origin to the initial estimate. This generation strategy is guided by prior calibration results, making it suitable for relatively fixed sensor configurations with moderate extrinsic variation. By positioning virtual cameras along the vector from the LiDAR origin to the initial estimated camera position, this approach can increase the overlapping field of view. However, it does not account for modality differences or feature density. In scenarios where LiDAR point clouds are considerably sparser than camera images, such as in KITTI and nuScenes, this method cannot effectively determine the optimal number of virtual cameras to balance calibration accuracy and computational efficiency. Furthermore, in robotic applications such as multi-robot joint perception, obtaining an initial estimate between sensors on different platforms is challenging. The method also depends on the accuracy of the initial guess; if the estimate is unavailable or inaccurate, the approach becomes ineffective. Given these limitations and the lack of an environmental perception mechanism, this strategy is impractical for environment-driven calibration.
We believe that these alternative methods for generating virtual cameras can be applied across a range of real-world scenarios. However, rather than seeking an elusive “best” method, it is more important to choose the most suitable configuration based on the specific application. We believe that concerning an algorithm module in an environment-driven framework, there is “no free lunch”. Users may balance the computational cost and calibration accuracy, selecting the approach that best suits their specific application to achieve more stable results. If the users are not sure which strategy to use, we strongly recommend using the default feature-density-based method, which provides a stable and generalizable approach for generating a sufficient number of virtual cameras.
VII Calibration Parameters
Calibration Parameters for Left RGB Camera
Calibration Parameters for Right RGB Camera
Examples of the data fusion result using our extrinsic calibration parameters in KITTI odometry are shown in the demo videos on our website.
VIII Details of Experimental Configuration
VIII-A Details of Comparison Methods
The details of the comparison methods are illustrated in Table IX. For algorithms that did not release their code, we directly reference the results published in their papers. Notably, RegNet [44], Borer et al.[24], and CalibDNN [27] employ different experimental datasets and evaluation processes, which limits the potential for a fully quantitative comparison. Furthermore, these methods do not specify which frames from the KITTI dataset were used for training and testing. Therefore, we rely on the published results to compare these approaches to others on the KITTI odometry 00 sequence.
@@ VIII-A Details of Comparison Methods
VIII-B Evaluation Metrics
To comprehensively evaluate the performance of LCEC approaches, we follow the previous works [25, 4, 28, 32], using the magnitude of Euler angle error and the magnitude of the translation error, with the following expression:
to quantify the calibration errors. In (25), and represent the estimated and ground-truth Euler angle vectors, computed from the rotation matrices and , respectively. Similarly, and denote the estimated and ground-truth translation vectors from LiDAR to camera, respectively. is the translation from LiDAR pose to camera pose when (1) is used to depict the point translation.
VIII-C Data Preparation
@@
We have conducted extensive experiments on KITTI odometry, KITTI360, nuScenes, MIAS-LCEC-TF70, and MIAS-LCEC-TF360. Now we provide our data preparation details.
@@
KITTI odometry: KITTI odometry is a large-scale public dataset recorded using a vehicle equipped with two RGB cameras Point Grey Flea 2 and one LiDAR of type Velodyne HDL-64E. Sensor data is captured at 10 Hz. We utilize the first 10 sequences (00-09) for evaluation. The dataset statistics of KITTI odometry are presented in Table X. It can be observed that there are a total of 24500 pairs of point clouds and images in KITTI odometry. Note that most of the SoTA comparison methods are slow to complete calibration with a single frame. For the comparison algorithms that provided source code, we performed calibration every ten frames (except for 2 frames in the 04 sequence), enhancing the efficiency of the experiments without compromising the validity of the evaluation.
KITTI-360: The KITTI-360 dataset is a large-scale autonomous driving dataset designed to advance research in various areas such as semantic scene understanding, 3D reconstruction, and motion forecasting. We utilize the images captured from the front stereo cameras and the point clouds scanned by the Velodyne HDL-64E LiDAR. We use the first 3000 pairs of LiDAR point clouds and camera images in the 00 sequence to test the performance of each component of EdO-LCEC. Similarly, we sampled the raw datasets every 10 frames.
nuScenes: The LiDAR sensor used in nuScenes is a 32-line spinning LiDAR (Velodyne HDL32E), which is considerably sparser than the 64-line LiDARs employed in KITTI and KITTI-360, thereby making the calibration task more challenging. To the best of our knowledge, few calibration methods have been evaluated on nuScenes. Following the official dataset split, we employ the nuScenes SDK to generate image-point cloud pairs from the 150 testing scenes in the v1.0-test set to construct our evaluation data. Since each scene in nuScenes contains only a limited number of sensor samples, we sample one out of every ten scenes to ensure the relative completeness of the data sequence within each scene.
MIAS-LCEC-TF70: MIAS-LCEC-TF70 is a diverse and challenging dataset that contains 60 pairs of 4D point clouds (including spatial coordinates with intensity data) and RGB images, collected using a Livox Mid-70 LiDAR and a MindVision SUA202GC camera, from a variety of challenging indoor and outdoor environments. We compare our method with SoTA approaches on the total six subsets to comprehensively evaluate the algorithm’s performance.
MIAS-LCEC-TF360: MIAS-LCEC-TF360 contains 12 pairs of 4D point clouds and RGB images, collected using a Livox Mid-360 LiDAR and a MindVision SUA202GC camera from both indoor and outdoor environments. Livox Mid-360 LiDAR produces a sparser point cloud compared to that generated by the Livox Mid-70 LiDAR. The significant difference in the FoV between this type of LiDAR and the camera results in only a small overlap in the collected data. Consequently, this dataset is particularly well-suited for evaluating the adaptability of algorithms to challenging scenarios.
Notably, an initial guess is recommended if the LiDAR and camera have a large difference in position or direction. For example, in the nuScenes dataset, the front camera and the LiDAR are installed in totally different directions with no shared field of view. We configure an initial offset on the translation and rotation to ensure that the initial mis-calibration rotation error is within 2 Euler degrees along each axis and the total translation error is within 0.5m.